Question 1
A company is aiming at raising $50 million debt for a new investment project. The company estimates that market investors would require 4.5% rate, whilst banks would charge 4.85%. The term of the debt is assumed to be 10 years. Considering that the company prefers to issue a 5% coupon bond (2.5% semi-annual), find:
1. The bond price at issue
2. The amount of debt issued (given the amount that the company wishes to collect)
3. The bond price one year after the issue if the new required rate is 4.8%
4. The amount of fees (in dollar and in % of the debt issued) that would make indifferent for the company issuing a corporate bond vs. taking up a bank loan.
Answer #1

The price of bond :

Ci MV (1 + YTM 4 (1+YTM)

or same can be calculate by use of excel function PV()

Answer a) Given Data as

Face Value $     1,000
Redemption Value as % of Par       100.00
Annual Coupon Rate 5.00%
Annual Required Return 4.50%
Semi annual required return 2.25%
Half yearly coupon (5/2=2.5%) $25
Maturity Time 10 years
Number of payment (10*2) 20
Value of Bond as % of Par       103.91%
Value of Bond in Dollars $ 1,039.91 PV(2.25%,20,-25,-1000)

Answer 2) Amount of debt issued = Required Fund / Value of bond = 50 million / 103.91% = $ 48.12 Million.

Answer 3) After 1 year ,

Face Value $     1,000
Redemption Value as % of Par       100.00
Annual Coupon Rate 5.00%
Annual Required Return 4.8%
Semi annual required return 2.4%
Half yearly coupon (5/2=2.5%) $25
Maturity Time 9 years
Number of payment (9*2) 18
Value of Bond as % of Par       101.45%
Value of Bond in Dollars $ 1,014.48 PV(2.4%,18,-25,-1000)

Answer 4) To make indifferent for the company issuing a corporate bond vs. taking up a bank loan.

Cost of bond should be equal to 4.85%.

Price of Bond at effective yield of 4.85% = PV(2.425%,20,-25,-1000) = $ 1011.78 or 101.18%

Cost of flotation per bond = 1039.91-1011.78 = $ 28.13

103.91%-101.18%= 2.73%